OpenStudy (anonymous):

If f(x) = −4x − 2, find f(−3)

OpenStudy (mayankdevnani):

f(-3)=-4(-3)-2 f(-3)=12-2 f(-3)=10

OpenStudy (mayankdevnani):

@brianhunter ok

OpenStudy (anonymous):

thanks

OpenStudy (mayankdevnani):

welcome